OpenStudy (anonymous):

Help Please! The number 19 to the power of start fraction five over three end fraction end power can be written in the form root of order cap a of 19 to the power of cap b end power. What is the value of B?

OpenStudy (phi):

\[ 19^\frac{5}{3} \] that can also be written this way \[ \left(19^5 \right)^\frac{1}{3} \] as you know , the exponent 1/3 means the same thing as the cube root

OpenStudy (anonymous):

my options are 19 3 5 5/3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

do you know that root of order cap a of 19 to the power of cap b end power. means \[ \sqrt[a]{19^b} \] ?